OpenStudy (anonymous):

The Ward family is redecorating several rooms of their house. Angelique wants to have a rug that is 9 feet long and 7 feet wide in her bedroom. The rug will cover the whole floor except a border that is x feet wide. The area of her room is 167 square feet. Find the width of the border, x. Round the answer to the nearest tenth of a foot.

OpenStudy (dumbcow):

 the area of border is: 9x + 9x+7x+7x+x^2 +x^2 + x^2+x^2 = 32x + 4x^2 --> 4x^2 +32x = 167-(9*7) = 104 --> x^2 +8x - 26 = 0 use quadratic formula to solve for x